# Semi-off-Policy Reinforcement Learning for Vision-Language Slow-thinking Reasoning

### v1

@@ 4.1 Semi-off-policy Sampling
Specifically, given a visual training dataset denoted as $\mathcal{D}_{\text{train}}=\{(I_{i},\bm{x}_{q_{i}},a_{i})|i=1,2,\cdots,|\mathcal{D}_{\text{train}}|\}$, where $I_{i}$ is the $i$-th image, $\bm{x}_{q_{i}}$ is the $i$-th query, $a_{i}$ is the $i$-th golden answer to $\bm{x}_{q_{i}}$ and ||$\cdot$|| ⋅ | calculates the size of set, the LVLM first performs visual understanding in an on-policy manner. For each image $I_{i}$, we first extract its visual tokens $\bm{v}_{i}=V(I_{i})$ using the pretrained visual encoder $V$ inside the LVLM. Then, prompt engineering [56] making the LVLM try its best to capture spatial layouts, semantic relationships among objects, fine-grained visual details and so on (Details in Figure A2 of Appendix E). Then, we can obtain a conditional distribution of image description $\bm{c}_{i}\sim\pi_{\theta}(\cdot|\bm{x}_{d},\bm{v}_{i})$.
@@ 6.3 Ablation Study
We assess the impact of weight freezing and the effectiveness of our proposed reward and evaluate them on MMMU, MathVista, MathVision and OlympiadBench. Further experiments on warm-up, response style, and caption as training data are detailed in Appendix D.
@@ 7 Conclusion and Future Work
Despite these advances, SOPHIA still struggles with long-range visual dependencies and fine-grained recognition in complex scenes. Also, performance may suffer when facing underrepresented domains or noisy samples. Issues such as hallucinated visual content, redundant reasoning, and the binary nature of reward signals still persist. Future work will explore stronger visual encoders, adaptive curriculum learning, and more robust data augmentation to further improve generalization, efficiency, and reasoning robustness across broader domains.
@@ Appendix A Theoretical Analysis
Alternatively, assume $\bar{\pi}(\bm{y}|\hat{\bm{x}}_{q})>0$. We then need to demonstrate that $R(\bm{x}_{q},\bm{v},\bm{y})=0$. If the specific tuple $(\bm{x}_{q},\bm{v},\bm{y})$ (for which we assume $\bar{\pi}(\bm{y}|\hat{\bm{x}}_{q})>0$) is indeed a sample in $\mathcal{D}$. Then, because $\bm{y}\notin\mathcal{Y}$, by the definition of $\mathcal{Y}$, this sample $(\bm{x}_{q},\bm{v},\bm{y})$ cannot satisfy $R(\bm{x}_{q},\bm{v},\bm{y})=1$. If it did, $\bm{y}$ would be in $\mathcal{Y}$, which contradicts our premise. Since the reward $R(\cdot)$ can only be 00 or $1$, it follows that $R(\bm{x}_{q},\bm{v},\bm{y})=0$.
@@ B.1 SOPHIA Setup
1:  Inputs: Training dataset $\mathcal{D}_{\text{train}}$, policy model (LVLM) $\pi_{\theta}$, reasoning language model $\bar{\pi}$, number of rollouts per image $K$, number of rollouts per caption $N$.
@@
8:        Sample $N$ reasoning trajectories $\{\bm{y}^{(1)},\bm{y}^{(2)},\cdots,\bm{y}^{(N)}\}$ where $\bm{y}^{(n)}\sim\bar{\pi}(\cdot|\hat{\bm{x}}_{q_{i}})$
@@ Appendix D Additional Experiment on SOPHIA
Warm up. To investigate the impact of the warm-up on the performance, we set up an experimental group including “SOPHIA w/o warm-up” and “SOPHIA w/o warm-up”. The policy initialization in the experimental group is directly InternVL2.5-38B (Instruct) and InternVL2.5-8B (Instruct) without a warm-up phase. The control group includes models trained following the standard procedure. The experimental results are shown in Table A1. We observe that models with warm-up consistently outperform those without warm-up. This is likely due to the emphasis of warm-up data on visual description and grounding. Such training enhances the captioning ability of model, thereby improving the quality of trajectories collected by the behavior model.
+SOPHIA 66.33 49.65 74.70 45.30 63.13 47.96 35.94 49.22 54.03
InternVL2.5-8B 48.22 28.61 64.30 26.78 39.14 21.02 20.61 12.90 32.70
+SOPHIA 57.00 32.02 61.10 30.84 46.51 22.66 22.00 18.53 36.33
Table A1: Ablation study of warm-up on InternVL2.5-38B and InternVL2.5-8B. A bold entry indicate best performance compared to others under the same base model.
Response Style. For the behavior policy, we evaluate different slow-thinking reasoning styles and their impact on SOPHIA performance. As shown in Fig.A1, the QwQ Preview style presents the full reasoning process without a summary or <think>…</think> tags, while the DeepSeek-R1 style is more decisive and uses these tags to separate reasoning from summary. Experiments are conducted on InternVL2.5-38B and no additional general visual or textual data are used. Results are shown in TableA2. We observe no significant differences across response styles, indicating that the focus at trajectories should be on the correctness and logical consistency of trajectories rather than style or format. In practice, we combine both response styles: QwQ Preview is used for simpler questions, while DeepSeek-R1 handles more complex ones.
@@
Keep-$N$. In our trajectory selection, we prioritize the shortest valid trajectory, which effectively implements a Keep-1 strategy [^1] Due to the length of trajectories and the use of sampling, duplicates are virtually impossible. A Keep-$N$ strategy selects the top-$N$ shortest trajectories for each image-query pair, as long as they are correct and their caption rewards exceed the threshold. This approach allows up to $N$ high-quality trajectories to be used for policy updates. This strategy may enhance learning in visual scenarios, and we evaluate its effectiveness on InternVL2.5-38B. Here, no additional general visual or textual data are introduced. Results are presented in Table A5. We find that the Keep-$N$ strategy can improve model performance to some extent, e.g., on OlympiadBench and MathVision. But the gains are marginal relative to the overall model abilities and come at a higher computational cost. Therefore, considering the trade-off between performance and efficiency, we adopt the Keep-1 strategy.
@@
Table A5: Pass@1 accuracy of SOPHIA on InternVL 2.5 38B in different $N$ ($N=1,2,8$).
@@ Appendix F Limitations
Despite the effectiveness of SOPHIA in enhancing visual slow-thinking reasoning, several limitations remain. First, LVLMs still struggle to retain long-range visual dependencies, especially in multi-hop or complex reasoning tasks, which can lead to partial or fragmented understanding of intricate visual scenes. Second, the current visual encoder lacks sufficient fine-grained recognition capability for complex scenes, which restricts the fidelity of visual perception and, consequently, the precision of downstream reasoning. Third, although the reward design in SOPHIA effectively filters out flawed trajectories, persistent issues such as hallucination and reasoning redundancy are not fully resolved, as evidenced by occasional overthinking or repetitive outputs in challenging benchmarks. Addressing these challenges in future work could involve integrating stronger visual grounding techniques, adaptive trajectory pruning, and multi-stage curriculum learning to further improve long-range visual reasoning and visual fidelity.

### v2

@@ 4.1 Semi-off-policy Sampling
Specifically, given a visual training dataset denoted as $\mathcal{D}_{\text{train}}=\{(I_{i},\bm{x}_{q_{i}},a_{i})|i=1,2,\cdots,|\mathcal{D}_{\text{train}}|\}$, where $I_{i}$ is the $i$-th image, $\bm{x}_{q_{i}}$ is the $i$-th query, $a_{i}$ is the $i$-th golden answer to $\bm{x}_{q_{i}}$ and $|\cdot|$ calculates the size of set, the LVLM first performs visual understanding in an on-policy manner. For each image $I_{i}$, we first extract its visual tokens $\bm{v}_{i}=V(I_{i})$ using the pretrained visual encoder $V$ inside the LVLM. Then, prompt engineering [56] making the LVLM try its best to capture spatial layouts, semantic relationships among objects, fine-grained visual details and so on (Details in Figure A2 of Appendix E). Then, we can obtain a conditional distribution of image description $\bm{c}_{i}\sim\pi_{\theta}(\cdot|\bm{x}_{d},\bm{v}_{i})$.
@@
We assume that the generated $\bm{c}_{i}$ can effectively encode the visual cues required for reasoning. Prior studies show that, in LLaVA-style LVLMs, adapter layers project visual embeddings into text-embedding subspaces, aligning continuous visual features with their textual representations [57]. Hence, visual understanding (whether expressed as embeddings or tokens) serves as the conditioning context for autoregressive reasoning. When this context captures spatial layouts, object relations and fine-grained details accurately, the reasoning becomes more coherent and logically consistent.
@@ 6.3 Ablation Study
We assess the impact of weight freezing, and the effectiveness of our proposed reward and warm-up. Further experiments on response style, teacher models, and caption as training data are detailed in Appendix D.
@@
Warm up. To investigate the impact of the warm-up on the performance, we set up an experimental group including “SOPHIA w/o warm-up” and “SOPHIA w/o warm-up”. The policy initialization in the experimental group is directly InternVL2.5-38B (Instruct) and InternVL2.5-8B (Instruct) without a warm-up phase. The control group includes models trained following the standard procedure. The experimental results are shown in Table 6. We observe that models with warm-up consistently outperform those without warm-up. This is likely due to the emphasis of warm-up data on visual description and grounding. Such training enhances the captioning ability of model, thereby improving the quality of trajectories collected by the behavior model.
+SOPHIA 66.33 49.65 74.70 45.30 63.13 47.96 35.94 49.22 54.03
InternVL2.5-8B 48.22 28.61 64.30 26.78 39.14 21.02 20.61 12.90 32.70
+SOPHIA 57.00 32.02 61.10 30.84 46.51 22.66 22.00 18.53 36.33
Table 6: Ablation study of warm-up on InternVL2.5-38B and InternVL2.5-8B. A bold entry indicate best performance compared to others under the same base model.
@@ 6.4 Analysis of Training Data
6.5 Qualitative Analysis on Hallucinations and Logical Inconsistencies
Method Visual Hallucination Logic Inconsistency
InternVL3.0-38B 38% 54%
InternVL3.0-38B+SOPHIA 16% 44%
Table 9: Qualitative analysis of SOPHIA.
We conduct manual evaluation on 50 randomly sampled cases from MMMU Pro before and after SOPHIA training, covering diverse disciplines to assess both visual hallucination and logical inconsistency. We report the proportion of responses with visual hallucinations and logic inconsistency. As shown in Table 9, SOPHIA significantly reduces hallucinations and logic error, consistent with improvements in overall correctness, which supports our motivation that on-policy visual understanding training can enhance the visual understanding of the model while reducing hallucinations.
@@ 7 Conclusion and Future Work
Despite these advances, SOPHIA still struggles with long-range visual dependencies and fine-grained or low-level recognition in complex scenes. Also, performance may suffer when facing underrepresented domains or noisy samples. Issues such as hallucinated visual content, redundant reasoning, and the binary nature of reward signals still persist. Future work will explore stronger visual encoders, adaptive curriculum learning, and more robust data augmentation to further improve generalization, efficiency, and reasoning robustness across broader domains.
Acknowledgements
We would like to thank the anonymous reviewers and area chairs for their comprehensive and constructive reviews. This work is supported by the Shanghai Artificial Intelligence Laboratory. The authors would like to thank Yiming Zhang, Chengqi Lyu and Jiangning Liu to their valuable suggestions and comments on this work.
@@ Appendix A Theoretical Analysis
Alternatively, assume $\bar{\pi}(\bm{y}|\hat{\bm{x}}_{q})>0$. We then need to demonstrate that $R(\bm{x}_{q},\bm{v},\bm{y})=0$. If the specific tuple $(\bm{x}_{q},\bm{v},\bm{y})$ (for which we assume $\bar{\pi}(\bm{y}|\hat{\bm{x}}_{q})>0$) is indeed a sample in $\mathcal{D}$. Then, because $\bm{y}\notin\mathcal{Y}$, by the definition of $\mathcal{Y}$, this sample $(\bm{x}_{q},\bm{v},\bm{y})$ cannot satisfy $R(\bm{x}_{q},\bm{v},\bm{y})=1$. If it did, $\bm{y}$ would be in $\mathcal{Y}$, which contradicts our premise. Since the reward $R(\cdot)$ can only be 0 or $1$, it follows that $R(\bm{x}_{q},\bm{v},\bm{y})=0$.
@@ B.1 SOPHIA Setup
1:Inputs: Training dataset $\mathcal{D}_{\text{train}}$, policy model (LVLM) $\pi_{\theta}$, reasoning language model $\bar{\pi}$, number of rollouts per image $K$, number of rollouts per caption NN.
@@
8:   Sample NN reasoning trajectories $\{\bm{y}^{(1)},\bm{y}^{(2)},\cdots,\bm{y}^{(N)}\}$ where $\bm{y}^{(n)}\sim\bar{\pi}(\cdot|\hat{\bm{x}}_{q_{i}})$
@@ B.2 GRPO Steup
B.3 Computational Costs
During the full training process, the computational cost of training the InternVL3.0-38B model serves as the baseline. Under the same data scale, the proposed GRPO-based training consumes roughly three-quarters of that baseline’s total computation. For inference, the SOPHIA-trained model follows the same algorithmic procedure as the base LVLM (i.e., next-token prediction). However, due to its slow-thinking and step-by-step reasoning design, it requires a longer reasoning sequence to reach the final output. Empirically, the inference process of SOPHIA is about 3-4 times slower than that of the base LVLM on standard reasoning benchmarks. This moderate increase aligns with the “test-time scaling” paradigm in reasoning models
@@ Appendix D Additional Experiment on SOPHIA
Response Style and Teacher Model. For the behavior policy, we evaluate different slow-thinking reasoning styles and their impact on SOPHIA performance. As shown in Fig.A1, the QwQ Preview style presents the full reasoning process without a summary or <think>…</think> tags, while the DeepSeek-R1 style is more decisive and uses these tags to separate reasoning from summary. Experiments are conducted on InternVL2.5-38B and no additional general visual or textual data are used. Results are shown in Table A1. We observe no significant differences between two different models in terms of quality (DeepSeek-R1 > QwQ-32B-Preview) and size (658B > 32B), indicating that the focus at trajectories should be on the correctness and logical consistency of trajectories rather than style or format, as well as SOPHIA can achieve strong performance without an extremely powerful teacher model. In practice, we combine both response styles: QwQ Preview is used for simpler questions, while DeepSeek-R1 handles more complex ones.
@@
Keep-NN. In our trajectory selection, we prioritize the shortest valid trajectory, which effectively implements a Keep-1 strategy [^1] Due to the length of trajectories and the use of sampling, duplicates are virtually impossible. A Keep-$N$ strategy selects the top-$N$ shortest trajectories for each image-query pair, as long as they are correct and their caption rewards exceed the threshold. This approach allows up to $N$ high-quality trajectories to be used for policy updates. This strategy may enhance learning in visual scenarios, and we evaluate its effectiveness on InternVL2.5-38B. Here, no additional general visual or textual data are introduced. Results are presented in Table A4. We find that the Keep-$N$ strategy can improve model performance to some extent, e.g., on OlympiadBench and MathVision. But the gains are marginal relative to the overall model abilities and come at a higher computational cost. Therefore, considering the trade-off between performance and efficiency, we adopt the Keep-1 strategy.
@@
Table A4: Pass@1 accuracy of SOPHIA on InternVL 2.5 38B in different NN ($N=1,2,8$).
@@ Appendix F Limitations
Despite the effectiveness of SOPHIA in enhancing visual slow-thinking reasoning, several limitations remain. First, LVLMs still struggle to retain long-range visual dependencies, especially in multi-hop or complex reasoning tasks, which can lead to partial or fragmented understanding of intricate visual scenes. Second, the current visual encoder lacks sufficient fine-grained recognition capability for complex scenes. We acknowledge that some low-level features, such as gradients and pixels, are not captured by semantic representations, since they are largely independent of semantics, which restricts the fidelity of visual perception and, consequently, the precision of downstream reasoning tasks that rely heavily on such features. Third, although the reward design in SOPHIA effectively filters out flawed trajectories, persistent issues such as hallucination and reasoning redundancy are not fully resolved, as evidenced by occasional overthinking or repetitive outputs in challenging benchmarks. Addressing these challenges in future work could involve integrating stronger visual grounding techniques, adaptive trajectory pruning, and multi-stage curriculum learning to further improve long-range visual reasoning and visual fidelity.